FEL Upgrade Project Monthly Report May 2002


By the end of the month more than 70% of the support stands for electron beam transport hardware have been installed in the FEL vault including the mounts for the IR wiggler and the second and third cryomodules. Three of the "girder assemblies" which contain recently released quadrupole magnets have been fully assembled and released for installation in the vault.

During a site visit to the manufacturer for the large 180 degree dipole magnets for the recirculations of the FEL Upgrade we witnessed a successful measurement of the flatness of the pole faces: they have been machined to a flatness of <0.0005 inches over the full 3 meter beam path.

A critical final test run for ion implantation of the Upgrade gun cathode support tube was successfully completed.


On May 1-3, the FEL Team hosted our 4th semiannual review of the FEL Upgrade project for our contract monitors in DOE, ONR and the AFRL. The review went well based on the preliminary feedback given to the FEL team and lb management t the close-out of the review. A formal report by the review committee was received in June.

The CRADA between JLab and AES, which was approved by DOE, was signed the early part of May by Jefferson Lab and forwarded to AES.

Since CRADA numbers and dollars seems to be on the increase, the process by which such agreements are drafted, reviewed, and finalized needs to be re-examined to see if there are ways to streamline the process without sacrificing requirements. A proposal has been drafted and submitted to Tech Transfer Manager for comment and additional streamlining suggestions prior to submitting to members of the Directors Council.

The loan agreement between Cornell and Jefferson Lab for the loan of Cornell’s undulator to Jefferson Lab for use on the UV FEL Upgrade was signed and forwarded to Cornell.

We prepared abstracts for 6 presentations/posters that were submitted to the Program Committee today (May 10th) for the 2002 International FEL Conference that will be held in September 9-13 in Chicago. These abstracts were forwarded to the ONR and AFRL program offices.

G. Neil attended a review of DOD’s Joint Technology Office funded programs at Albuquerque on May 6-8.

On Tuesday, May 14, George Neil and Fred Dylla presented a summary of the FEL development program to the DARPA panel on electric ship systems held at Syntek Systems in Arlington, VA. The presentation was well received.

On Friday, May 17th, several members of the FEL group and the Accelerator Division participated in a jointly sponsored workshop at the University of Virginia on potential applications of the FEL and future lights sources for analyzing biostructures and nanomaterials.

The FEL project management team (Dylla and Neil) attended a meeting of the FEL technical Working Group (FEL-TAWG) on Wednesday May 22 sponsored by the DoD Joint Technology Office. The TAWG is being re-constituted under a new chair Michael Dietchman, from ONR, who is taking over after the retirement of Eli Zimet in January. Many of the FEL experts that have been serving on the FEL Advisory Committee for the FEL Upgrade project have been invited to be members of the TAWG. The purpose of this working group is to advise the JTO on FEL R&D.

On Thursday, May 23 we attended a workshop in Washington on directed energy activities sponsored by the Directed Energy Professional Society (DEPS). The Navy's interest and ownership of the FEL development was presented by RADM Mathis and Cmd. Roger McGinnis of NAVSEA. The JLab FEL program was given good press at this meeting by both the above speakers and by the secretary of last year's Defense Science Board panel on high energy laser development.

F. Dylla presented a talk on applications of the JLab FEL to surface analysis and surface processing to the AVS sponsored Topical Conference on Surface Analysis held on May 21-23 at Vanderbilt University.

FEL project management and JLab management hosted a meeting of our Maritime Technical Advisory Committee (MTAC) on May 28-29. MTAC has advised the lab on FEL development for Navy interests since the FEL program was formally kicked-off in 1996. MTAC is chaired by Admiral Al Baciocco, and includes Admiral Ted Parker, Admiral Ken Malley and the former director of the Naval Research Labs, Alan Berman, as members. We have always appreciated the wise and experienced council offered by MTAC members and were again grateful for their cogent and frank input. We received a formal report from MTAC in June.

On Wednesday May 29th, George Neil attended a meeting of the Navy R&D teams who are participating in the laser materials damage (lethality) studies funded by the DOD Joint Technology Office. Planning continued for the next phase of lethality measurements to be done using the 10 kW Upgrade along with our partners from NRL and MIT-Lincoln Labs.

Project Cost Performance:

The Phase 1 FEL Upgrade project budget for the period June 1, 2000 to September 30, 2001 is $9,029k. The first phase of the project was fully obligated as of the end of September. The total Phase 1 project cost to date-actual costs plus commitments is $9,029k as planned for Phase 1 completion.

The Phase 2 FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002. Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k. Work for Phase 2 proceeded to plan during this reporting period. Actual charges of $259k were accrued for May for total accrued costs of $3,277k to date for Phase 2. Work scheduled for May was $571k for a total of $2,680k for Phase 2 to date; work performed for May was $357k for a total of $2,795k for Phase 2 to date, which results in 60% scheduled vs. 63% performed for the Phase 2 project. At present the cost variance for Phase 2 is -$482k and the schedule variance is $115k. We expect to meet the planned cost at completion of Phase 2 effort. The program is presently on track to begin commissioning efforts in September 2002.

The UV FEL project was approved for $2,836k for a performance period from September 2001 to September 30, 2002. Actual charges of $506k were accrued through May for a total accrued costs of $1,339k since project start. Work scheduled for May was $459k for a total of $1,415k for the UV project to date, work performed for May was $387k for a total of $1,482k for the UV project to date, which results in 50% scheduled vs. 52% performed for the UV project. At present the cost variance for the UV project is $144k and the schedule variance is $68k. The program is on track to meet the statement of work requirements on schedule and within the available budget.

WBS 3 (Beam Physics):

We spent some time this month analyzing and reviewing the quads. The quads demonstrably can perform to spec but give different results for B(I) curves depending on the timing of the excitation (as expected). Unfortunately, we can't time the hysteresis loop in the same way in both the vault and on the test stand (different power supply systems), so the measured B (I) curve doesn't necessarily predict the achieved (in the vault) values to within spec (unfortunately, not as expected).

Discussions with MEG, CASA, controls, and EEG group members clarified the issue, and a plan to provide measurement capability matching operational practice was worked up.

WBS 4 (Injector):

The support tube polishing was completed and is ready for implant. A step forward in the work with W&M on the implant process for the support tube was achieved. Reducing the operating pressure and a thorough cleaning of the plasma source chamber solved the problem of arcs observed in the first test tube. The second test tube shows no evidence of arcs after two runs. A third test run showed some arcing that was traced to dust that had accumulated at the interface of the appendage chamber holding the support tube when the chamber was flushed with gas. Since the gas can’t be flushed through the appendage, the flushing procedure will be replaced with a chamber vacuum cleaning, a HEPA filter will be mounted above the chamber, and gas line filters and CO2 dusters. The shield door will be implanted next with the ball cathode to follow.

Continued working on PARMELA simulations to meet the e-beam parameters at the match point by finding the proper quadrupoles settings.

Worked on the designs for the SF6 tank modifications and the HV resistor.

We held a design review for the lightbox for the gun and signed off the drawings this month.

The gun chamber was cleaned and mounted back on the stand; it is now ready for installation in the vault. Design work continued on the SF6 can modifications

Gun HVPS - The delivery of the SF6 tank was moved back out to 6/16/02 by the vendor. They had a problem with some flanges they purchased for the tank. The first conditioning resistor was sprayed with the resistive ink and air dried. The incremental weight gain and film thickness checked out.

WBS 6 (RF):

Completed HV wiring of both CPSs for 100 kW klystrons with the exception of wiring to klystron carts.

The plumbing has been installed upstairs for the klystron carts on the LCW system. We have moved the 100 kW klystron cart into position for unit #3 due to the relocation of the magnet box power supply into FL01 zone.


Quarter HVPS - Completed low level wiring of unit 3 and started testing Heater Voltage circuits. Heater Voltage board for unit 4 is finished, but not yet installed. The PLC code is not expected until 3rd week of June. New dual 18 volt power supplies were ordered for the klystron heater monitor circuits. The circuits were tested and found to be linear over the range of interest. The low level wiring is finished for one unit and nearly finished for the other. A 50 kW klystron is being connected as a test load for the HVPS. The PLC programming and any new wiring is the next step before testing can begin.

Quarter Klystrons - The 2nd and 3rd klystron carts are finished at the Machine Shop. They will be delivered and the klystrons installed early next month.

WBS 8 (Instrumentation):

1) Cable tray straightening was completed

2) 1st article 8 channel stepper motor chassis was turned over to Al for system debug (good results)

3) IOC FEL3 was upgraded to Epics r3132j0 and the timer card software and hardware was changed out to provide a one-board system. Preliminary testing is good.

4) CANbus interface for the beamviewers was turned over to Jeff Wilson with a completion date around May 30.

5) Electrical plumbing for the new dipole box supply has begun.

6) The rest of the 3" beam viewers have had the interface box converted, tested and turned over to the vacuum group(c/o J. Heckman) for installation.

7) The water supply (feed/return, manifolds and valves) for the super trim magnet racks have been identified and some parts ordered.

8) The preexisting beamviewer cameras CCD's have been tested and repaired.

9) Re-terminated trim magnet feeds in Junction Box 1and 2.

10) Fabrication of four 4 channel programmable digital pulse generator boards were completed

DC Power: The east arc box power supply was connected to the AC lines, and feed for the new upgraded west arc supply is being worked. The old pi bends were run from a single power supply with a 100 amp feed, the new arc supplies are fed from 300 amp each. An additional transformer and panel will be installed for the 12 - 100amp/50 volt power supplies. The transformer/panel will use the 100-amp feed from the old pi bends. The 208-volt power could have been scrounged from the length of the gallery but being able to lock out all supplies from a single location is much safer and saves both labor and materials from using many power panels. A new water manifold is also being installed for the 100 amp'ers

An effort was made at cleaning up noise from the magnet test stand and to identify sources of drift. We have made progress and testing of quads resumed.

The lens mirror covers for the beamviewers were painted and ready for installation onto the frames. Progress was made on the cable pull for the Ion Pumps. The new smoke detectors were interlocked into the LSS. The PLCs for the MPS, CAMAC crate and modules for the BLMs have been removed from the racks along with a clean up of the cables.

FL06B01 (in the drive laser clean room) has been upgraded to include a VME crate running a new ioc (iocfel10) for the new drive laser pulse control system. This effort included a reorganization of existing hardware and a re-working of the cabling in the rack. Work continues on the Inventory Control System and significant improvements have been posted to the server. The first attempt at a check print for the BLM vme card (F0151) is being reviewed and appended.

The quadrupole girder assembly has moved into full swing. They are being assembled and delivered at 1 or 2 a day, depending on complexity. There are a few parts that are still trickling, like shielded pump drops, but this should not hold up assembly. Improvements continue to be made on the magnet test stand; the PCI based motor driver code is being done by Mike Necaise. This will begin the transition out of CAMAC to an entirely PC/LabView based test stand. This work had been scheduled for some time by Tom Hiatt and crew, but with the quads now moving on the decision was made to begin the conversion. Many thanks to the magnet test guys and the vacuum assembly crews!

Fabrication nearly complete for 3 of the 4 new OMS chassis. One OMS chassis has been installed in the control room for further electrical testing and the software installed for compatibility. The mechanical drawing for the BLM photo multiplier assembly has been revised and awaiting sign off. Design upgrade for the raster power supplies has begun. Maintenance for the LeCroy HV mainframe has been completed, tested and re-installed in the rack. Continuing cable clean up in the quarter cryo region. Website maintenance and upgrades are continuing. The schematic for the beam viewer Canbus interface has been finished and PCB layout has begun.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• The first eight coils have completed their surface repair. They potted their tenth coil. The white spots have gone away to be replaced by two other puzzling phenomena.

• After being removed from the mold and looking beautifully translucent, the outer surface of the epoxy starts to take on a haze which is theorized as a separation from the glass fibers in zones of the ground wrap between conductors. According to MEI, this happens to coils that are potted with Nadic Methyl Anhydride (NMA) hardener. They usually take care of it with an Acrylic or Silicone sprayed-on coating. (This is why you see painted coils, so the imperfections are hidden) MEI has conducted some tests of treated vs. untreated glass tape with no conclusive results. They are going to try treating a portion of the mold with an alternate, non-silicone based mold release that they can trust and see if the coil in that portion doesn’t have the haze.

• The second phenomenon is a residual of a brown oily residue on the surface of the coil that is easily removed with xylene. We are in consultation with Jeff Hubrig, our epoxy consultant this residue is from the potting process and is benign

• Upper and lower core’s plates are in final machining with projection that first article assembly will be at the start of next month.

• MEI made the fixtures for gluing the Purcell gap shims in place and they have some trial shim pieces and a trial core piece that they will do prototype gluing with before they will do the real gluing.

Injector Dipoles (DU/DV)

• At Wang NMR, they are working on potting the coils.

• At Master Machine, core pieces remain in final machining. All hardware and tooling is completed, first article will be assembled next month.

Bend/Reverse Bend Dipoles (GX, GQ)

• Process Equipment Co. (PECo) continued manufacturing the cores and parts of the GQ-GX dipoles. They completed heat-treating of GQ core material and started final machining. GX material is in heat-treating process.

• Wang NMR continues working on the winding fixture for the coils. They are setting up a winding line separate from the GY coil for these coils.

• AES completed the run on GX dipoles at high energy with a larger shim only on the IR portion of the large face. Dave Douglas is evaluating it.

• AES drew up the shim system for the use of these magnets at the highest energy. The drawings were checked and sent to PECo for incorporation into their production.

180 Degree Dipoles (GY)

• Wang NMR completed winding the first GY coil, reworked the winding fixture and started winding the second coil.

• Bosma Machine complete assembly of the first article GY Core. We inspected it, finding that the gap height was within the .0005 inch (a quarter of a human hair) uniformity required. It is beautiful. After some additional painting, they shipped it to Wang NMR for assembly.

• Wang NMR is working on the potting fixture.


3 inch quad (QX)

• We decided to re measure all the QX quads to the new CEBAF, full range ("bang-bang") hysteresis protocol. However, further tests showed that the power supply we use and the trim rack power supplies do not have the same characteristics. As the month wore on, our decision to re measure become more complex. Though the magnet quality is proven, the measurements have been perpetually fogged by drift and noise, Tom Hiatt asked Kevin Jordan to look at their whole measurement system. The look by an electronic/electrical expert has revealed a legacy of processes and instruments for measuring magnets that are archaic (from the original CEBAF Commissioning days). They do not compare to present day methods used to run hysteresis loops and set magnets in the CEBAF accelerator. As a result, test protocol was simplified to 17 minutes rather than 45 and will exactly match the magnet setting method. Kevin tackled the noise and drift but was not able to stop it.

• We found that the electric current set point reliability of the Trim Card Power Supplies is about ±3 parts in 1000, well below David Douglas’ specification for magnetic field set ability of 1 part per 1000. Grist for getting better trim supplies. After some down time on the magnet test stand, the full production retesting of the QX quads started and 6 QX quads were released for installation.

Trim Quad (QT)

• Milhous Control of Virginia continues making the remaining nine magnets. We found that the first article had a leak in the cooling circuit from a post manufacturing clean out of a water fitting.

• Vertical Corrector (GC) As a quick analog computation, (Using the device itself) we tested the quad at low current to assess whether by powering coils appropriately, it would be able to have the superposed corrector built into it. Results were such that we will try it on our magnetic model.

Sextupole (SF)

• DULY Research continues to work on completing drawings of the magnet. They are concentrating on designing the splice and cooling water blocks on the coils so that when brazed, the small passage is not blocked and so that the manifolds and connections between the six coils will be compact but serviceable. Details of SF are complete except for stand mounting. They are coordinating with JLAB designer to complete the mounting. Preliminary stand/support design is done.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• The drawings for ARC Chambers were checked, corrected and signed off. The procurement documents for the ARC Chambers are in composition.

• Special region drawings for Arcs, Wiggler, Injector dipoles, Optical Chicane and Light Box were given to the Survey and Alignment Group with preliminary graphics but approved stand and cartridge positions following a final check. The Alignment Group laid out the position of the latest group of stands and a number of the stands in the first arc (They came in this week) were placed in position.

• Additional design personnel were assigned to the recirculation region in front of the Cryomodule.

• Design of the recirculation region in front of the Cryomodule is nearing completion.

• Design of the optical chicane region with its chamber is about half done.

• The only region still not completed or in process is the recirculation dump.

WBS 10 (Wiggler):

The wiggler was installed onto its stands this week and is now ready for alignment and diagnostic installation

WBS 11 (Optics):

We met with members of the Survey and Alignment group to determine if the number and size of access ports in the optical cavity vacuum vessels is adequate. Agreement was reached on a strawman alignment plan, and this is incorporated into the drawings. An assembly and test fixture was partially "mocked up" in User Lab 2, and based on the layout, sketches created and submitted to the machine shop for fabrication. The electrical feedthroughs for the vacuum vessels have been ordered. We received the leadscrews and nuts for the optical cavities and the assembly/test stands.

We have a scheme for translating the flexible cooling water hoses and electrical conduit, and have ordered parts to test it. We are designing the backplane cooling for the mirror and turning cassettes. We have a budgetary quote for the insertable mirrors (which are water cooled) and the bid seems high so we are checking other sources.

While we achieve acceptable performance from the feedback loop on the OCMMS (active stabilization) electronics, we would like it to be ~ 5X better. Plans for the new circuit were made; smaller and higher resolution position sensitive detectors were ordered. Work continues on the drive laser pulse compressor. We installed a spectrometer in the drive laser room to measure the bandwidth, and thus to determine the pulsewidth. Sketches to provide mounting fixtures to In-braze mirrors and work with Ga:In eutectic were submitted to the machine shop for fabrication. The vendor of our ultrafast laser system informed us that the Ti:sapphire oscillator should ship today.

We attended meetings reviewed plans for laser purchases, and created a preliminary design for an external resonant cavity for the FEL (aka the pulse stacking cavity). This is primarily funded by a DIRUP grant to the College of William and Mary (M. Kelley - P.I.)

Our active mirror stabilization effort has involved a great deal of work by an Old Dominion Univ. EE student, Mr. Chris Behre. Chris prepared and gave a presentation before a panel of three engineers, and prepared a paper as well. His work was awarded a 2nd place out of 17 entries. We're all proud of Chris, and congratulate him on his graduation, which took place on the third weekend of May.

We held informal reviews with optics group staff on the optical cavity vacuum vessels. The few suggestions were incorporated into the design, and the team shifted to producing detailed drawings. We hope to release these next week.

Our vendor for the optical cavity internals requested input on how to fabricate the yaw plate. We are meeting with our vacuum group and experts within the FEL team. Fabrication of the mounts that provides translation for these internal components are in process at the Lab Machine Shop

The same outside vendor is completing the OCMMS fabrication. They report they will complete one assembly this month (about 1 week late). Another assembly is delayed due to a slip in delivery of calcium fluoride vacuum viewports.

Mirror mounting tests with sapphire substrates are proceeding. A sapphire mirror mounted using In:Ag braze in A Ni-plated holder was temperature-cycled (warmed to ~ 30 deg. C above ambient, then cooled) with no change in its mounting integrity. A 2" ZnSe mirror's edge was sputter-coated with Ni and will be used for similar mounting tests. Nickel test fixtures that mimic the mounting arrangement on the Upgrade outcoupler holder (but with various gaps) will be used next week to determine the optimum gap required to obtain a uniformly dispensed bead of Ga:In eutectic, another candidate heat conduction medium.

The fabrication of the first article 50kW optical beam dump is nearly complete. A second design, employing "foam metal" (a metal mesh) in the cooling channels to promote turbulent flow and thus increase heat transfer coefficients has been awarded to an outside vendor and is being fabricated.

Our one setup for laser beam profiling is becoming oversubscribed, so we ordered additional hardware and software. The equipment arrived this week and we'll install it for immediate use to relieve the bottleneck. Software that will be used for interpreting laser pulse shapes was installed and tested using simulated data. Initial measurements of the drive laser output after using fiber-grating pulse compression indicate that the pulse has been compressed from a value of ~ 60 ps to a few ps, as expected from reports in the literature.

The chiller used to temperature-stabilize the ultrafast laser system arrived, joining the pump laser. We expect the oscillator this month, which will complete the acquisition of this important piece of hardware. An analysis to determine eye hazards from this laser system was conducted, and appropriate eyewear selected and ordered. Wiring in UL 1 to provide

safety interlocking for the new laser was also installed this week.

An additional review of the design of the optical component for the THz radiation monitor resulted in a change to return the electro-optical crystal to a position in air. This simplified the design. We hope to complete this design effort in the next two weeks.

Detailing the drawings for the optical cavity vacuum vessels has been somewhat more time-consuming than originally projected, so the plan is to release preliminary drawings to the fabrication scheduling team next week, and final drawings the week after. We met and discussed the request for fabrication advice by the vendor of the optical cavity internals. We determined that the SOW contains sufficient guidance, and signed off on the vendor's plan to machine both sides of all plates.

Fabrication the first article 50kW optical beam dump is complete, except for brazing. This will be done within the next two weeks. The second "foam metal" (a metal mesh) design is also complete, but requires the metal mesh before it can be brazed. Vendors to create the absorber on the copper were identified and will be contacted to get pricing and delivery.

Our additional laser beam profiler arrived and was installed. We met with the software engineers to iterate plans for getting the new IR beam profiler output into EPICS.

The THz radiation monitor mechanical design was reviewed with respect to minimizing wakefields. Solutions were proposed (adding a trapdoor) and this is being incorporated into the design.

We are looking at ways to induce chirp on the electron beam (and thus the FEL output) in such a way that the electron beam can still be energy recovered. This will be needed when we wish to compress the pulsewidth of the FEL.


Aerospace Corp. April 2002 Progress Report

During this reporting period the design team began to formalize the design of the Jefferson Laboratory micromachining station. A visit was conducted to Aerotech Corporation (Pittsburgh, PA), one of the premier vendors for computer controlled-motion. Aerotech routinely supplies hardware to the semiconductor industry and has developed numerous XYZ motion control systems for laser micromachining applications. Although, Aerotech has made "high speed" hardware for laser micromachining, the laser repetition rates typically are KHz and not the expected MHz repetition rate of the JLab UV FEL. We discussed our desire to utilize the MHz repetition rate of the JLab FEL and requested Aerotech to review their control systems and the new FireWireTM driven software. Our goal was to see if it can meet the needs for very high speed processing at JLab. While visiting their applications laboratory, we conducted several trials with existing hardware to test the ability for abrupt motion maneuvers at velocities near 1m/sec. The results showed that vibration, in the acoustical domain is minimal at these very high speeds but we agreed that a laser vibrometer measurement will be needed to measure the spectrum.

An oral presentation was prepared for the midyear Review. The presentation included a high level design of the envisioned deliverable system (at end of CY 03 but subject to CY03 funding) and a program mitigation strategy given the CY02 expected funding shortfall of >$700K.


Aerospace Corp. May 2002 Progress Report

A presentation was given to the UV/IR FEL Program Review Committee for the Midyear Review. The presentation showed that the tasks, deliverables and budget are on target for completion in CY02. The presentation also highlighted the major funding shortfall (>$700K) expected from the FY02 funds and how this shortfall will impact the development of the micromachining station and overall project. It is believed that if this shortfall can be recovered with funds from the FY03 budget and both the FY02 and FY03 funds could be delivered together in early CY03, then little impact is expected on the delivery of the JLab UV FEL Micromachining Station (end of CY03).

Aerotech Corporation engineers visited the Aerospace design team in Los Angeles to discuss engineering aspects of their very high speed XYZ motion control system. Critical aspects of their software and hardware control were discussed. Aerospace prepared and sent to Aerotech examples of typical motion sequences that will be run at the factory under controlled conditions. Additional discussions were held regarding the development and nuances of the Post-Processor (Post) software. The Post is a necessary piece of software that translates commands generated by the tool-path software to vendor specific hardware (G-code) language. After evaluation of various vendors, Aerospace has chosen MasterCAM as the tool-path software. MasterCAM has the market share in computer assisted manufacturing (CAM) software. This CAM software is used by most CNC machine tools and student versions are also available. Aerospace will deliver a professional license seat with the JLab micromachining station to allow users a simple means for generating tool-paths from their CAD drawings.

Aerospace also hosted a visit by Kevin Jordan of JLab. We presented to JLab our current ideas and approaches, as regarding overall system control, motion control, laser safety, remote access control and our desire to utilize the JLab MHz repetition rate effectively. In return, Aerospace received a presentation on the existing IR FEL control systems that are in place, the types of software and hardware control system available and how the Aerospace deliverable machine can be "connected" to the overall JLab FEL system. We agreed to explore the JLab desire to make uniform the drivers and circuitry that control pico-motors, cameras and A/Ds for non real-time tasks. In our case these would be motors that move mirrors and other optics into position prior to initiating three-dimensional high-speed processing.